|-(-13)|-|8|=

-7(-5+2)(4+9)+7=
-14-6-9-20-15+19+6=
pls help me with this problem
thank you very much

Answer by CubeyThePenguin(3113)

|-(-13)| - |8| = |13| - |8| = 13 - 8 = 5

-7(-5 + 2)(4 + 9) + 7
= -7(-3)(13) + 7
= 273 + 7
= 280

-14 - 6 - 9 - 20 - 15 + 19 + 6 = -39